40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1034     Entry time: Wed Oct 8 19:17:55 2008
Author: Yoichi 
Type: Configuration 
Category: PSL 
Subject: Laser power is slowly recovering 
This afternoon we (rich, steve, yoichi) shutdown the laser for the DC-DC converter installation.
(we decided not to do so. Detail will be posted soon.)
After we turned on the laser again,the laser power has been recovering but very slowly.
At the time of writing, the laser power is 2.6W (MOPA_AMPMON).
I think it is because the chiller temparature has not yet settled down (it went up to 25C and slowly coming down, now at 22C).
It will take some hours until the power fully comes back.
Right now I confirmed that at least the MC locks.
ELOG V3.1.3-